4. Find the ratio e : f, if 5e2 – 13ef – 6f2 = 0 where e and f ≠ 0.


1
Expert's answer
2021-06-25T07:10:48-0400
"5e^2-13ef-6f^2=0"

Divide both sides by "f^2\not=0"


"\dfrac{5e^2}{f^2}-\dfrac{13ef}{f^2}-\dfrac{6f^2}{f^2}=0"

"5(\dfrac{e}{f})^2-13(\dfrac{e}{f})-6=0"

Solve the quadratic equation for "\dfrac{e}{f}"


"D=(-13)^2-4(5)(-6)=289"

"\dfrac{e}{f}=\dfrac{13\pm\sqrt{289}}{2(5)}"

"\dfrac{e}{f}=\dfrac{13-17}{10}=-0.4"

or


"\dfrac{e}{f}=\dfrac{13+17}{10}=3"

"e: d=-2:5" or "e:f=3:1"